Levothyroxine sodium is a synthetic form of the hormone thyroxine (T4), which increases the metabolic rate. A typical adult maintenance dose is 1.7 micrograms/day for every kilogram of body mass. For a patient who weighs 194 pounds, how many 25-microgram tablets of levothyroxine sodium must be taken each day?

While drugs such as levothyroxine sodium are used to treat hypothyroid conditions, drugs such as methimazole are given to control hyperthyroidism. In severe cases, an adult loading dose can be as large as 30 mg per administration, with two administrations per day. If 84 tablets are prescribed for a two-week methimazole regimen, how many milligrams of methimazole are present in each tablet?

Calcium chloride is used to treat disorders arising from low levels of calcium in the blood plasma. A typical adult dose is 500 mg/day administered from 1 to 3 days. This dose of medicine is delivered via a slow intravenous injection. If the injection takes 5 minutes to deliver the drug, what is the dosage rate in milligrams per minute?

Another treatment for low calcium levels is ergocalciferol (vitamin D2). An average adult dose is around 2.5 mg/day. If taken in the liquid form, which has a concentration of 0.2 mg/ml, how many milliliters of the drug solution would one daily dose contain?
Answer #1
  1. A typical adult maintenance dose of Levothyroxine sodium= 1.7 micrograms/day for every kilogram of body mass.

For a patient who weighs 194 pounds=88.18 kg

Required amount=88.18X1.7 micrograms=149.90mg=150 mg

How many 25-microgram tablets of levothyroxine sodium must be taken each day?

=150/25=6 tablets.

  1. Adult loading dose can be=30 mg per administration, with two administrations per day= means 60mg per day

Means for two weeks required amount=60X14=840 mg

Given that 84 tablets are prescribed for a two-week methimazole regimen= 840 mg

How many milligrams of methimazole are present in each tablet?

=840/84

=10mg of methimazole is present in each tablet.

  1. A typical adult dose of calcium chloride= 500 mg/day

administered from 1 to 3 days.

This dose of medicine is delivered via a slow intravenous injection. If the injection takes 5 minutes to deliver the drug.

The dosage rate in milligrams per minute=500/5

=100 mg/min.

  1. An average adult dose of ergocalciferol (vitamin D2) is= 2.5 mg/day.

If taken in the liquid form, which has a concentration of 0.2 mg/ml,

How many milliliters of the drug solution would one daily dose contain?

0.2 mg supplied from 1ml

then for 1mg required ml= 5ml

Then 2.5 mg will be supplied from=2.5 X5

12.5 ml of the drug solution would one daily dose contain